OpenStudy (anonymous):

wilL someone pLease heLp me . . . Identify the classifications that apply to the real number. √12 a.)natural number, whole number, integer, rational number, terminating decimal b.)irrational c.)natural number, whole number, integer, rational number d.)rational number, terminating decimal

OpenStudy (anonymous):

The real number in question is sqrt(12)?

OpenStudy (anonymous):

sqrt(12) = sqrt(3*4) = 2sqrt(3)

OpenStudy (anonymous):

sqrt(3) is irrational.

OpenStudy (anonymous):

and we know that the product of a rational and irrational (nonzero) is always irrational.

OpenStudy (anonymous):

idont understand . . .

OpenStudy (anonymous):

First did I understand the question right?

OpenStudy (anonymous):

\[\sqrt{12} = \sqrt{3*4} = 2*\sqrt{3}\]

OpenStudy (anonymous):

is that okay? sqrt(4)=2 so we can "pull" that out.

OpenStudy (anonymous):

yes ! im tryinq to fiqure out what appLys to that # .

OpenStudy (stacey):

If the number does not have a straight forward square root, then it is irrational. Sqrt(2) and sqrt(3) are irrational. On the other hand sqrt(0.09) = 0.3, so that is rational.